OpenStudy (anonymous):

Solve a2 - 81 = 0 by factoring. Select all possible answers. -9 0 No solution Infinite solutions 9

OpenStudy (anonymous):

do you know this rule ? a^2 - b^2 = (a-b)*(a+b). If no, learn it, you will use it many times :) Now, you can write a^2 - 81 = 0 as (a-9)*(a+9) = 0 so you have a=9 and a=-9

OpenStudy (anonymous):

I don't know that rule.

OpenStudy (anonymous):

for clarification: 81 =9^2

OpenStudy (anonymous):

ok then you should ask yourself: what number should i choose and take its square and that is equal to 81 ? Apparently 9 and -9 because 9^2 =81 and (-9)^2 = 81, too

OpenStudy (anonymous):

oh..